OpenStudy (anonymous):

PLEASE HELP ME !!!!! How do I turn this cos function into a sin function : 19 cos(0.104t)21

OpenStudy (tkhunny):

You're going to have to enter it better than that. sin(a) = cos(pi/2 - a)

OpenStudy (anonymous):

sorry 19 cos(0.104 t) + 21

OpenStudy (anonymous):

So will it be 19 sin(0.104(t-pi)) +21?

OpenStudy (tkhunny):

\(\cos(0.104t) = \sin\left(\dfrac{\pi}{2} - 0.104t\right)\)

OpenStudy (anonymous):

so when you graph both of them they should look the same right ?

OpenStudy (tkhunny):

And indeed they do!

OpenStudy (anonymous):

they do not look a like

OpenStudy (anonymous):

l don't understand

OpenStudy (tkhunny):

Then you did not graph them both correctly. This is no good: sin(0.104(t-pi)) This is right: sin(pi/2 - 0.104t)

OpenStudy (anonymous):

Okay but where did you find pi/2 ?

OpenStudy (anonymous):

like is there a mathematical way to find that shift ?

OpenStudy (tkhunny):

Way back in your trigonometry class, you were forced to memorize and manipulate identity after identity after identity after identity. Are you remembering this with fondness? This is just one of those many identities. \(\sin(x) = \cos(\pi/2 - x)\) It's just an identity. You can use \(\sin^{2}(x) + \cos^{2}(x) = 1\) if you like. I just thought I'd pick an easier one and save some trouble.

OpenStudy (anonymous):

okay so I insert the cos and I will get that

OpenStudy (tkhunny):

Pretty hard to confirm via rumor. Show me what you are planning.

OpenStudy (anonymous):

sin^2 +cos(0.104) ^2 +1

OpenStudy (tkhunny):

?? That makes no sense at all. You have \(\sin^{2}(x) + \cos^{2}(x) = 1\) This is an identity. As long as those arguments are the same, the sum of the two functions will be 1. \(\sin^{2}(2) + \cos^{2}(2) = 1\) \(\sin^{2}(\sqrt{2}) + \cos^{2}(\sqrt{2}) = 1\) \(\sin^{2}(Fred) + \cos^{2}(Fred) = 1\) \(\sin^{2}(0.104t) + \cos^{2}(0.104t) = 1\) If you have \(\cos(0.104t)\), you must solve the identity for that value or function. \(\cos^{2}(0.104t) = 1 - \sin^{2}(0.104t)\) \(\cos(0.104t) = \sqrt{1 - \sin^{2}(0.104t)}\) -- Although, you have to be a little careful with that, since we just messed with the Domain. I highly recommend the original substitution.

OpenStudy (anonymous):

Oh I see thank you !

OpenStudy (tkhunny):

?? That makes no sense at all. You have \(\sin^{2}(x) + \cos^{2}(x) = 1\) This is an identity. As long as those arguments are the same, the sum of the two functions will be 1. \(\sin^{2}(2) + \cos^{2}(2) = 1\) \(\sin^{2}(\sqrt{2}) + \cos^{2}(\sqrt{2}) = 1\) \(\sin^{2}(Fred) + \cos^{2}(Fred) = 1\) \(\sin^{2}(0.104t) + \cos^{2}(0.104t) = 1\) If you have \(\cos(0.104t)\), you must solve the identity for that value or function. \(\cos^{2}(0.104t) = 1 - \sin^{2}(0.104t)\) \(\cos(0.104t) = \sqrt{1 - \sin^{2}(0.104t)}\) -- Although, you have to be a little careful with that, since we just messed with the Domain. I highly recommend the original substitution.